OpenStudy (anonymous):

If a sprinkler waters 1 over 15 of a lawn in 1 over 3 hour, how much time will it take to water the entire lawn? 5 hours 15 hours 18 hours 45 hours

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

a??

OpenStudy (michele_laino):

I'm trying...

OpenStudy (michele_laino):

please, do you know what is the entire lawn?

OpenStudy (anonymous):

15

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

please note that you rate is: \[\frac{ 1+1/15 }{ 1+1/3 }=\frac{ 16/15 }{ 4/3 }=\frac{ 16 }{ 15 }*\frac{ 3 }{ 4 }=\] \[=\frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

so then is it a??

OpenStudy (michele_laino):

sorry, your rate is: \[r=\frac{ 1/15 }{ 1/3 }=\frac{ 1 }{ 5 }\] then necessary time is: \[time=\frac{ 15 }{ 1/5 }=75\]

OpenStudy (anonymous):

so then the answer is...

OpenStudy (michele_laino):

are you sure that entire lawn is 15?

OpenStudy (michele_laino):

I think is A., namely the first option

OpenStudy (michele_laino):

@77777jeannie77777

OpenStudy (anonymous):

ok... thx

OpenStudy (anonymous):

Sam measured the height of a window frame as 5.23 feet, but the actual height was 6 feet. What is the percentage of error in Sam's measurement? 6.79% 7.79% 11.23% 12.83%

OpenStudy (michele_laino):

explanation: since sprinkler uses 1/3 to water 1/15 of lawn, he have to use 15 times 1/3 in order to water th entire lawn, namely 15*(1/3)=5 hours

OpenStudy (anonymous):

what do you think @Michele_Laino

OpenStudy (michele_laino):

oops, ...he has to use....

OpenStudy (michele_laino):

I think is the first option, namely t=5 hours

OpenStudy (anonymous):

@Michele_Laino